OpenStudy (anonymous):

Write and solve the system of equations below. Let a represent adults and let c represent children. 1,200 tickets were sold to the spring carnival. Adult tickets cost $6.00 and children's tickets cost $3.50. A total of $6,450 was collected. How many tickets of each kind were sold?

OpenStudy (texaschic101):

a + c = 1200 ---> a = 1200 - c 6a + 3.5c = 6450 now sub 1200 - c in for a in the other equation 6(1200 - c) + 3.5c = 6450 -- distribute through the parenthesis 7200 - 6c + 3.5c = 6450 -- subtract 7200 from both sides -6c + 3.5c = 6450 - 7200 -- combine like terms -2.5c = -750 -- divide by -2.5 c = 300 a + c = 1200 a + 300 = 1200 a = 1200 - 300 a = 900 check.. 6a + 3.5c = 6450 6(900) + 3.5(300) = 6450 5400 + 1050 = 6450 6450 = 6450 (correct) adult tickets sold (a) = 900 child tickets sold (c) = 300